Please answer this question :)

1) Write out any sequence of DNA, 21 bases in length a) Assumung this 21 base DNA fragment is a primer, calculate the TM of this fragment b )What approximate temperature does each G or C base contribute to the primer fragment c) what temperature does each A or T base contribute? d) What temperature should the annealing temperature of this fragment be?
Answer #1

1) ATCCAGGACAAGGAGGGCATA

a) TM of a primer = 4 x (no. of G and C) + 2 x (no. of A and T)

Therefore, TM of the above primer = (4x11) + (2x10) = 44+20 = 64oC

b) Each G or C base contributes for 4oC to the primer fragment

c) Each A or T base contributes for 2oC to the primer fragment

d) The annealing temperature of this fragment should be 64 plus or minus 5oC